The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
2/12

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2 / 12 = (1× 2) / (6 × 2)
=1 / 6

Thanks